OpenStudy (anonymous):

if there are 0.036 grams of HCL dissolved in one liter of solution, what i the pH?

OpenStudy (anonymous):

First, you will find the amount of HCl in moles:\[n _{HCl}=\frac{ m }{ M }=\frac{ 0.036 }{ 36.45 }\approx 10^{-3}(mol)\]Next, you need to calculate the concentration of H+: {H+]=n/V. Finally, to get the pH, you might want to use the formula: pH = -log([H+])